# Event-Customized Image Generation

### v1

@@ Abstract
Customized Image Generation, generating customized images with user-specified concepts, has raised significant attention due to its creativity and novelty. With impressive progress achieved in subject customization, some pioneer works further explored the customization of action and interaction beyond entity (i.e., human, animal, and object) appearance. However, these approaches only focus on basic actions and interactions between two entities, and their effects are limited by insufficient “exactly same” reference images. To extend customized image generation to more complex scenes for general real-world applications, we propose a new task: event-customized image generation. Given a single reference image, we define the “event” as all specific actions, poses, relations, or interactions between different entities in the scene. This task aims at accurately capturing the complex event and generating customized images with various target entities. To solve this task, we proposed a novel training-free event customization method: FreeEvent. Specifically, FreeEvent introduces two extra paths alongside the general diffusion denoising process: 1) Entity switching path: it applies cross-attention guidance and regulation for target entity generation. 2) Event transferring path: it injects the spatial feature and self-attention maps from the reference image to the target image for event generation. To further facilitate this new task, we collected two evaluation benchmarks: SWiG-Event and Real-Event. Extensive experiments and ablations have demonstrated the effectiveness of FreeEvent.
Figure 1: Customized Image Generation. (a) Generating customized images with given subjects in new contexts. (b) Generating customized images with co-existing basic action or interaction in given images. (c) Generating customized images for complex events with various target entities. Different colors and numbers show the associations between reference entities and their corresponding target prompts.
@@ 1 Introduction
Recently, large-scale pre-trained diffusion models (Dhariwal & Nichol, 2021; Nichol et al., 2021; Ramesh et al., 2022; Rombach et al., 2022; Saharia et al., 2022) have demonstrated remarkable success in generating diverse and photorealistic images from text prompts. Leveraging these unparalleled creative capabilities, a novel application — customized image generation (Gal et al., 2022; Ruiz et al., 2023; Chen et al., 2023) — has gained increasing attention for generating user-specified concepts. Significant progress has already been made in subject-customized image generation (Ye et al., 2023; Chen et al., 2024b). As shown in Figure 1(a), given a set of user-provided subject images, existing methods can accurately capture the unique appearance features of each subject (e.g., corgi) with a special identifier token, enabling creative rendering in new and diverse scenarios. Moreover, they can seamlessly integrate multiple subjects into cohesive compositions, preserving their distinctive characteristics while adapting them to novel contexts.
@@
Beyond the appearance of different entities (i.e., humans, animals, and objects) in the images, pioneering approaches have been developed to customize the user-specified actions (Huang et al., 2024), interactive relations (Huang et al., 2023) and poses (Jia et al., 2024) between the entities. As shown in Figure 1(b), these methods attempt to capture the single-entity action (e.g., handstand) or interactions (e.g., back to back) between two entities that co-exist in the given reference images and transfer them to the synthesis of action- or interaction-specific images with new entities.
@@
However, for real-world scenes that typically involve multiple entities with more complex interactions (e.g., Figure 1(c), row three: three humans are discussing in front of a computer with different poses), these works (Huang et al., 2023; 2024; Jia et al., 2024) still face notable limitations. 1) Simplified Customization. Current action customization (Huang et al., 2024) focuses solely on the basic actions of a single person. Similarly, interaction customizations (Huang et al., 2023; Jia et al., 2024) are limited to basic interactive relations or poses between just two entities. There is a lack of exploration into more complex and diverse actions or interactions that involve multiple humans, animals, and objects. Additionally, while these methods typically perform well when generating images with the same type of entity (e.g., all monkeys or all cats), they struggle when faced with more diverse and complex entities and their combinations. These narrow focuses and limitation on entity generation have strictly limited their abilities to customize more complex and diverse scenes with creative content. 2) Insufficient Data. To capture specific actions or interactions, existing methods (Huang et al., 2023; 2024; Jia et al., 2024) tend to represent them by learning corresponding identifier tokens, which can be further used for generating new images. However, for each action, or interaction, these training-based processes typically require a set of reference images (e.g., 10 images) paired with corresponding textual descriptions across different entities. Unfortunately, each action or interaction is highly unique and distinctive, i.e., gathering images that depict the exact same action or interaction is challenging. As shown in Figure 1(b), there are still significant differences in the same action (e.g., handstand) between different reference images, which thus compromises the accuracy of learned tokens, leading to inconsistencies in action between generated images. This insufficient data issue for identical action or interaction has severely limited the practicality and generalizability of these methods.
@@
Moreover, as a pioneering effort in this direction, we also collected two evaluation benchmarks from the existing dataset (i.e., SWiG (Pratt et al., 2020) and HICO-DET (Chao et al., 2015)) and the internet for event-customized image generation, dubbed SWiG-Event and Real-Event, respectively. Both benchmarks include reference images featuring diverse events and entities, along with manually crafted target prompts. Extensive experiments demonstrate that our approach achieves state-of-the-art performance, enabling more complex and creative customization with enhanced practicality and generalizability.
In summary, we make several contributions in this paper: 1) We propose the novel event-customized image generation task, which extends customized image generation to more complex scenes in real-world applications. 2) We propose FreeEvent, the first training-free method for event customization, which can be further combined with subject customization methods for more creative and generalizable customizations. 3) We collect two evaluation benchmarks for event-customized image generation, and our FreeEvent achieves outstanding performance compared with existing methods.
@@ 2 Related Work
Text-to-Image Diffusion Generation. Diffusion models (Ho et al., 2020; Nichol & Dhariwal, 2021; Song et al., 2020) have emerged as a leading approach for image synthesis. The text-to-image diffusion models (Nichol et al., 2021; Ramesh et al., 2022; Saharia et al., 2022) further inject user-provided text descriptions into the diffusion process via pre-trained text encoders. After trained on large-scale text-image pairs, they have shown great success in text-to-image generation. Different from these models that operate the diffusion process on pixel space, the latent diffusion models (LDMs) (Rombach et al., 2022) propose to perform it on latent space with enhanced computational efficiency. Besides, existing works (Hertz et al., 2022; Tumanyan et al., 2023; Cao et al., 2023; Alaluf et al., 2024) have discovered the spatial feature and attention maps in LDMs contain localized semantic information of the image and the layout correspondence between textual conditions. As a result, these features and attention maps have been utilized to control the layout, structure, and appearance in text-to-image generation. This can be achieved either through a plug-and-play feature injection (Tumanyan et al., 2023; Xu et al., 2023; Lin et al., 2024) or by computing specific diffusion guidance (Epstein et al., 2023; Mo et al., 2024) for generation. In this paper, we utilize the pre-trained LDM Stable Diffusion (Rombach et al., 2022) as our base model.
@@
Subject Customization. This task aims to generate customized images of user-specified subjects. Current mainstream subject customization works mainly focus on 1) Single subject customization, including learning specific identifier tokens (Gal et al., 2022), finetuning the text-to-image diffusion model (Ruiz et al., 2023; 2024), introducing layer-wise learnable embeddings (Voynov et al., 2023) and training large-scale multimodal encoders (Gal et al., 2023; Li et al., 2024). 2) Multi-subject composition, including cross-attention modification (Tewel et al., 2023), constrained model fine-tuning (Kumari et al., 2023), layout guidance (Liu et al., 2023), and gradient fusion of each subject (Gu et al., 2024). In conclusion, these works are all tailored to capture the appearance of the entities in the image, without considering the customization of actions or poses.
@@
Action and Interaction Customization. They aim to generate customized images with co-existing actions or interactions in user-provided reference images. ReVersion (Huang et al., 2023) first proposes to customize specific interactive relations by optimizing the learnable relation tokens. ADI (Huang et al., 2024) makes progress in customizing specific actions for a single subject. And a following work (Jia et al., 2024) further extends it to learning interactive poses between two individuals. However, all these works only focus on simplified customization of some basic actions and interactions, and their effect is strictly limited by the insufficient data of reference images. In contrast, our proposed event customization only requires one reference image, and our training-free framework FreeEvent can achieve effective customization of complex events with various creative target entities.
@@ 3.1 Preliminary
Latent Diffusion Model. Generally, the LDMs include a pretrained autoencoder and a denoising network. Given an image $x$, the encoder $\mathcal{E}$ maps the image into the latent code $z_{0}=\mathcal{E}(x)$, where the forward process is applied to sample Guassian noise $\mathbf{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ to it to obtain $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ from time step $t\sim\left[1,T\right]$ with a predefined noise schedule $\bar{\alpha}$. While the backward process iteratively removes the added noise on $z_{t}$ to obtain $z_{0}$, and decodes it back to image with the decoder $x=\mathcal{D}(z_{0})$. Specifically, the diffusion model is trained by predicting the added noise $\mathbf{\epsilon}$ conditioned on time step $t$ and possible conditions like text prompt $\mathrm{P}$. The training objective is formulated as
@@
Diffusion Guidance. The diffusion guidance modifies the sampling process (Ho et al., 2020) with additional score functions to guide it with more specific controls like object layout (Xie et al., 2023; Mo et al., 2024) and attributes (Epstein et al., 2023; Bansal et al., 2023). We express it as
@@
3.2 Task Definition: Event-customized Image Generation
In this section, we first formally define the event-customized image generation task. Given a reference image $I^{R}$ involves $N$ reference entities $\mathrm{E}^{R}=\{R_{1},\dots,R_{N}\}$, we define the “event" as the specific actions and poses of each single reference entity, and the relations and interactions between different reference entities. Together we have the entity masks $\mathrm{M}=\{m_{1},\dots,m_{N}\}$, where $m_{i}$ is the mask of its corresponding entity $R_{i}$. The event-customized image generation task aims to capture the reference event, and further generate a target image $I^{G}$ under the same event but with diverse and novel target entities $\mathrm{E}^{G}=\{G_{1},\dots,G_{N}\}$ in the target prompt $\mathrm{P}=\{w_{0},\dots,w_{N}\}$, where $w_{i}$ is the description of the target entity $G_{i}$, and each target entity $G_{i}$ should keep the same action or pose with its corresponding reference entity $R_{i}$. As the example shown in Figure 2, given the reference image with four reference entities (e.g., three people and one object), the event-customization aims to capture the complex reference event and generate the target image with a novel combination of different target entities (e.g., skeleton, statue, monkey, book).
@@ 3.3 Approach
U-Net Architecture The Stable Diffusion (Rombach et al., 2022) utilizes the U-Net architecture (Ronneberger et al., 2015) for $\epsilon_{\theta}$, which contains an encoder and a decoder, where each consists of several basic encoder/decoder blocks, and each encoder/decoder block further contains several encoder/decoder layers. Specifically, as shown in Figure 3(a), each U-Net encoder/decoder layer consists of a residual module, a self-attention module, and a cross-attention module. For block $b$, layer $l$, and timestep $t$, the residual module produces the spatial feature of the image as $\mathbf{f}$. The self-attention module produces the self-attention map as $\mathbf{SA}=\mathrm{Softmax}(\frac{\mathbf{Q}_{s}\mathbf{K}_{s}^{\mathrm{T}}}{\sqrt{d}})$, where $\mathbf{Q}_{s}$ and $\mathbf{K}_{s}$ are query and key features projected from the visual features. For text-to-image generation, the cross-attention module further produces the cross-attention map between the text prompt $\mathrm{P}$ and the image as $\mathbf{CA}=\mathrm{Softmax}(\frac{\mathbf{Q}_{c}\mathbf{K}_{c}^{\mathrm{T}}}{\sqrt{d}})$, where $\mathbf{Q}_{c}$ is the query features projected from the visual features, and $\mathbf{K}_{c}$ is the key features projected from the textual embedding of $\mathrm{P}$.
Entity Switching Path. This path aims on generating target entities $\mathrm{E}^{G}=\{{G}_{1},\dots,G_{N}\}$ in $I^{G}$ by switching each refenrece entity ${R}_{i}$ to ${G}_{i}$ based on the target prompt $\mathrm{P}$ and reference entity masks $\mathrm{M}$. And the key is to ensure each target entity ${G}_{i}$ is generated at the same location as their corresponding reference entity ${R}_{i}$ and avoid the appearance leakage between different entities. Inspired by prior works (Hertz et al., 2022; Chen et al., 2024a) that utilize the cross-attention maps to control the layout of text-to-image generation, we apply the cross-attention guidance and regulation to achieve the entity switching.
As shown in Figure 2(a), at the timestep $t$ of the denoising process, we first obtain the latent for entity switching as $z^{A}_{t}=z^{G}_{t}$, we then input $z^{A}_{t}$ together with the target prompt $\mathrm{P}$ into the U-Net, and calculate the cross-attention maps as $\mathbf{{CA}}^{A}$. As shown in Figure 3(b), we then introduce an energy function to bias the cross-attention of each token $w_{i}$ as:
@@
where $\mathrm{{CA}}^{A}_{i}$ is the cross-attention map of token $w_{i}$. Optimizing this function encourages the cross-attention maps of each target entity ${G}_{i}$ to obtain higher values inside the corresponding area specified by $m_{i}$, and further guide the localized layout of each target entity. We calculate the gradient of this guidance via backpropagation to update the latent $z^{G}_{t}$:
@@
Event Transferring Path. This path aims to extract the specific reference event from the reference image $I^{R}$, including the action, pose, relation, or interactions between each reference entity, and transferring them to the target image $I^{G}$. Meanwhile, from the perspective of image spatial information, the event is essentially the structural, semantic layout, and shape details of the image. Thus, based on the observation that the spatial features and self-attention maps can be utilized to control the image layout and structure (Tumanyan et al., 2023; Xu et al., 2023; Lin et al., 2024), we perform spatial feature and self-attention map injection to achieve the event transferring.

### v2

@@ Abstract
Customized image generation has raised significant attention due to its creativity and novelty. With impressive progress achieved in subject customization, some pioneer works further explored the customization of action and interaction beyond entity (e.g., human and object) appearance. However, these approaches only focus on basic actions and interactions between two entities, and their effects are limited by insufficient “exactly same” reference images. To extend the customization to more complex scenes, we propose a new task: event-customized image generation. Given a single reference image, we define the “event” as all specific actions, poses, relations, or interactions between different entities in the scene. This task aims at accurately capturing the complex event and generating customized images with various target entities. To solve this task, we proposed a novel training-free method: FreeEvent. Specifically, FreeEvent introduces two extra paths alongside the general diffusion denoising process: 1) Entity switching path: it applies cross-attention guidance and regulation for target entity generation. 2) Event transferring path: it injects the spatial feature and self-attention maps from the reference image to the target image for event generation. We further collected two new evaluation benchmarks. Extensive experiments have demonstrated the effectiveness of FreeEvent.
Diffusion Model, Image Generation
Figure 1: Customized Image Generation. (a) Generating customized images with given subjects in new contexts. (b) Generating customized images with co-existing basic action or interaction in given images. (c) Generating customized images for complex events with various target entities. Different colors and numbers show associations between reference entities and corresponding target prompts.
@@ 1 Introduction
Recently, large-scale pre-trained diffusion models (dhariwal2021diffusion; nichol2022glide; ramesh2022hierarchical; rombach2022high; saharia2022photorealistic) have demonstrated remarkable success in generating diverse and photorealistic images from text prompts. Leveraging these unparalleled creative capabilities, a novel application — customized image generation (gal2023image; ruiz2023dreambooth; chendisenbooth) — has gained increasing attention for generating user-specified concepts. Significant progress has already been made in subject-customized image generation (ye2023ip; chen2024anydoor). As shown in Figure 1(a), given a set of user-provided subject images, existing methods can accurately capture the unique appearance features of each subject (e.g., corgi) with a special identifier token, enabling creative rendering in new and diverse scenarios. Moreover, they can seamlessly integrate multiple subjects into cohesive compositions, preserving distinctive characteristics while adapting them to novel contexts.
@@
Beyond the appearance of different entities (i.e., humans, animals, and objects) in the images, pioneering approaches have been developed to customize the user-specified actions (huang2024learning), interactive relations (huang2024reversion) and poses (jia2024customizing) between the entities. As shown in Figure 1(b), these methods attempt to capture the single-entity action (e.g., handstand) or interactions (e.g., back to back) between two entities that co-exist in the given reference images and transfer them to the synthesis of action- or interaction-specific images with new entities.
@@
However, for real-world scenes that typically involve multiple entities with more complex interactions (e.g., Figure 1(c), row 3: three humans are discussing in front of a computer with different poses), these works (huang2024reversion; huang2024learning; jia2024customizing) still face notable limitations. 1) Simplified Customization. Current action customization (huang2024learning) focuses solely on the basic actions of a single person. Similarly, interaction customizations (huang2024reversion; jia2024customizing) are limited to basic interactive relations or poses between just two entities. There is a lack of exploration into more complex and diverse actions or interactions that involve multiple humans, animals, and objects. Additionally, while these methods typically perform well when generating images with the same type of entity (e.g., all monkeys or all cats), they struggle when faced with more diverse and complex entities and their combinations. These narrow focuses and limitation on entity generation have strictly limited their abilities to customize more complex and diverse scenes with creative content. 2) Insufficient Data. To capture specific actions or interactions, existing methods (huang2024reversion; huang2024learning; jia2024customizing) tend to represent them by learning corresponding identifier tokens, which can be further used for generating new images. However, for each action, or interaction, these training-based processes typically require a set of reference images (e.g., 10 images) paired with corresponding textual descriptions across different entities. Unfortunately, each action or interaction is highly unique and distinctive, i.e., gathering images that depict the exact same action or interaction is challenging. As shown in Figure 1(b), there are still significant differences in the same action (e.g., handstand) between different reference images, which thus compromises the accuracy of learned tokens, leading to inconsistencies in action between generated images. This insufficient data issue for identical action or interaction has severely limited the practicality and generalizability of these methods.
@@
Moreover, as a pioneering effort in this direction, we also collected two evaluation benchmarks from the existing dataset (i.e., SWiG (pratt2020grounded) and HICO-DET (chao2015hico)) and the internet for event-customized image generation, dubbed SWiG-Event and Real-Event, respectively. Both benchmarks include reference images featuring diverse events and entities, along with manually crafted target prompts. Extensive experiments demonstrate that our approach achieves state-of-the-art performance, enabling more complex and creative customization with enhanced practicality and generalizability.
In summary, we make several contributions in this paper: 1) We propose the novel event-customized image generation task, which extends customized image generation to more complex scenes in real-world applications. 2) We propose FreeEvent, the first training-free method for event customization, which can be further combined with subject customization methods for more creative and generalizable customizations. 3) We collect two evaluation benchmarks for event-customized image generation, and FreeEvent achieves outstanding performance compared with existing methods.
@@ 2 Related Work
Text-to-Image Diffusion Generation. Diffusion models (ho2020denoising; nichol2021improved; song2021denoising) have emerged as a leading approach for image synthesis. The text-to-image diffusion models (nichol2022glide; ramesh2022hierarchical; saharia2022photorealistic) further inject user-provided text descriptions into the diffusion process via pre-trained text encoders. After trained on large-scale text-image pairs, they have shown great success in text-to-image generation. Different from these models that operate the diffusion process on pixel space, the latent diffusion models (LDMs) (rombach2022high) propose to perform it on latent space with enhanced computational efficiency. Besides, existing works (hertz2023prompt; tumanyan2023plug; cao2023masactrl; alaluf2024cross) have discovered the spatial feature and attention maps in LDMs contain localized semantic information of the image and the layout correspondence between textual conditions. As a result, these features and attention maps have been utilized to control the layout, structure, and appearance in text-to-image generation. This can be achieved either through a plug-and-play feature injection (tumanyan2023plug; xu2023infedit; lin2024ctrl) or by computing specific diffusion guidance (epstein2023diffusion; mo2024freecontrol) for generation. In this paper, we utilize the pre-trained LDM Stable Diffusion (rombach2022high) as our base model.
@@
Subject Customization. This task aims to generate customized images of user-specified subjects. Current mainstream subject customization works mainly focus on 1) Single subject customization, including learning specific identifier tokens (gal2023image), finetuning the text-to-image diffusion model (ruiz2023dreambooth; ruiz2024hyperdreambooth), introducing layer-wise learnable embeddings (voynov2023p+) and training large-scale multimodal encoders (gal2023encoder; li2024blip). 2) Multi-subject composition, including cross-attention modification (tewel2023key), constrained model fine-tuning (kumari2023multi), layout guidance (liu2023cones), and gradient fusion of each subject (gu2024mix). In conclusion, these works are tailored to capture the appearance of the entities in the image, without considering the customization of actions or poses.
@@
Action and Interaction Customization. They aim to generate customized images with co-existing actions or interactions in user-provided reference images. ReVersion (huang2024reversion) first proposes to customize specific interactive relations by optimizing the learnable relation tokens. ADI (huang2024learning) makes progress in customizing specific actions for a single subject. And a following work (jia2024customizing) further extends it to learning interactive poses between two individuals. However, all these works only focus on simplified customization of some basic actions and interactions, and their effect is strictly limited by the insufficient data of reference images. In contrast, our proposed event customization only requires one reference image, and our training-free framework can achieve effective customization of complex events with various creative target entities.
@@ 3.1 Preliminary
Latent Diffusion Model (LDM). Generally, LDMs include a pretrained autoencoder and a denoising network. Given an image $x$, the encoder $\mathcal{E}$ maps the image into the latent code $z_{0}=\mathcal{E}(x)$, where the forward process is applied to sample Guassian noise $\mathbf{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ to it to obtain $z_{t}=\sqrt{\bar{\alpha}_{t}}z_{0}+\sqrt{1-\bar{\alpha}_{t}}\epsilon$ from time step $t\sim\left[1,T\right]$ with a predefined noise schedule $\bar{\alpha}$. While the backward process iteratively removes the added noise on $z_{t}$ to obtain $z_{0}$, and decodes it back to image with the decoder $x=\mathcal{D}(z_{0})$. Specifically, the diffusion model is trained by predicting the added noise $\mathbf{\epsilon}$ conditioned on time step $t$ and possible conditions like text prompt $\mathrm{P}$. The training objective is formulated as:
@@
Diffusion Guidance. The diffusion guidance modifies the sampling process (ho2020denoising) with additional score functions to guide it with more specific controls like object layout (xie2023BoxDiff; mo2024freecontrol) and attributes (epstein2023diffusion; bansal2023universal). We express it as
@@
3.2 Task Definition: Event-customized Generation
In this section, we first formally define the event-customized image generation task. Given a reference image $I^{R}$ involves $N$ reference entities $\mathrm{E}^{R}=\{R_{1},\dots,R_{N}\}$, we define the “event” as the specific actions and poses of each single reference entity, and the relations and interactions between different reference entities. Together we have the entity masks $\mathrm{M}=\{m_{1},\dots,m_{N}\}$, where $m_{i}$ is the mask of its corresponding entity $R_{i}$. The event-customized image generation task aims to capture the reference event, and further generate a target image $I^{G}$ under the same event but with diverse and novel target entities $\mathrm{E}^{G}=\{G_{1},\dots,G_{N}\}$ in the target prompt $\mathrm{P}=\{w_{0},\dots,w_{N}\}$, where $w_{i}$ is the description of the target entity $G_{i}$, and each target entity $G_{i}$ should keep the same action or pose with its corresponding reference entity $R_{i}$. As the example shown in Figure 2, given the reference image with four reference entities (e.g., three people and one object), the event-customization aims to capture the complex reference event and generate the target image with a novel combination of different target entities (e.g., skeleton, statue, monkey, book).
@@ 3.3 Approach
U-Net Architecture The Stable Diffusion (rombach2022high) utilizes the U-Net architecture (ronneberger2015u) for $\epsilon_{\theta}$, which contains an encoder and a decoder, where each consists of several basic encoder/decoder blocks, and each encoder/decoder block further contains several encoder/decoder layers. Specifically, as shown in Figure 3(a), each U-Net encoder/decoder layer consists of a residual module, a self-attention module, and a cross-attention module. For block $b$, layer $l$, and timestep $t$, the residual module produces the spatial feature of the image as $\mathbf{f}$. The self-attention module produces the self-attention map as $\mathbf{SA}=\mathrm{Softmax}(\frac{\mathbf{Q}_{s}\mathbf{K}_{s}^{\mathrm{T}}}{\sqrt{d}})$, where $\mathbf{Q}_{s}$ and $\mathbf{K}_{s}$ are query and key features projected from the visual features. For text-to-image generation, the cross-attention module further produces the cross-attention map between the text prompt $\mathrm{P}$ and the image as $\mathbf{CA}=\mathrm{Softmax}(\frac{\mathbf{Q}_{c}\mathbf{K}_{c}^{\mathrm{T}}}{\sqrt{d}})$, where $\mathbf{Q}_{c}$ is the query features projected from the visual features, and $\mathbf{K}_{c}$ is the key features projected from the textual embedding of $\mathrm{P}$.
Entity Switching Path. This path aims on generating target entities $\mathrm{E}^{G}=\{{G}_{1},\dots,G_{N}\}$ in $I^{G}$ by switching each refenrece entity ${R}_{i}$ to ${G}_{i}$ based on the target prompt $\mathrm{P}$ and reference entity masks $\mathrm{M}$. And the key is to ensure each target entity ${G}_{i}$ is generated at the same location as their corresponding reference entity ${R}_{i}$ and avoid the appearance leakage between different entities. Inspired by prior works (hertz2023prompt; chen2024training) that utilize the cross-attention maps to control the layout of text-to-image generation, we apply the cross-attention guidance and regulation to achieve the entity switching.
As shown in Figure 2(a), at the timestep $t$ of the denoising process, we first obtain the latent for entity switching as $z^{A}_{t}=z^{G}_{t}$, we then input $z^{A}_{t}$ together with the target prompt $\mathrm{P}$ into the U-Net, and calculate the cross-attention maps as $\mathbf{{CA}}^{A}$. Then, we introduce an energy function to bias the cross-attention of each token $w_{i}$ as (cf., Figure 3(b)):
@@
where $\mathrm{{CA}}^{A}_{i}$ is the cross-attention map of token $w_{i}$. Optimizing this function encourages the cross-attention maps of each target entity ${G}_{i}$ to obtain higher values inside the corresponding area specified by $m_{i}$, which further guides the localized layout of each target entity. We calculate the gradient of this guidance via backpropagation to update latent $z^{G}_{t}$:
@@
Event Transferring Path. This path aims to extract the specific reference event from the reference image $I^{R}$, including the action, pose, relation, or interactions between each reference entity, and transferring them to the target image $I^{G}$. Meanwhile, from the perspective of image spatial information, the event is essentially the structural, semantic layout, and shape details of the image. Thus, based on the observation that the spatial features and self-attention maps can be utilized to control the image layout and structure (tumanyan2023plug; xu2023infedit; lin2024ctrl), we perform spatial feature and self-attention map injection to achieve the event transferring.
@@
[…]
ControlNet 10.64 26.12 36.82 10.66 23.98 31.28 0.6009 0.3714 0.2599 0.2198 70.45
MIGC 10.90 27.00 37.64 10.62 26.14 35.04 0.6456 0.3772 0.2467 0.2145 49.81
BoxDiff 8.60 22.48 32.08 5.58 14.52 19.42 0.5838 0.3135 0.2099 0.2153 68.49
FreeEvent 41.12 63.02 72.74 34.10 62.04 71.82 0.7044 0.6282 0.5116 0.2238 29.05
Table 1: Performance of our model and state-of-art conditional text-to-image generation models on SWiG-Event. For image retrieval, the R@k represents that among the top-k images with the highest similarity to the target image, its corresponding reference image is included. For verb detection, the T-K represents the top-k detection accuracy.